OpenStudy (anonymous):

solve the following summation: sigma(k=10 to 70)(7k-2)

OpenStudy (anonymous):

Sigma (7k-2) = Sigma (7k)-Sigma 2 Sigma (7k) = 7Sigma k = 7(10+20 + ... + 70) = 7*280 = 1960 Sigma 2 = 2*Sigma 1 = 2*(1+1+...+1)= 2*7 =14 Sigma (7k-2) = 1960 - 14 = 1946

OpenStudy (anonymous):

Your method is good but \[\sum_{10}^{70}\] means that you need to add all the numbers between 10 and 70 not just 10/20/30.../60/70 , so you missed quite a lot.

OpenStudy (anonymous):

Ok then you'll have to use the formula for Sigma k: Sn = (10+70)*60/2 that means the first term +the last term multiplied by the number of terms and divided by 2 For Sigma 1 = (1+1+...+1) = 60 Sigma (7k-2) = 7*2400 -2*60 = 16680

OpenStudy (anonymous):

still not 100% correct I'm afraid :-) How many numbers are there between 10 and 70 including 10 and 70?